OpenStudy (anonymous):

Simplify 6 x 3 divide 3 - 1 Answers: 2 3 5 6

OpenStudy (anonymous):

6*3=18/3=6-1=5 so C *Your welcome

OpenStudy (anonymous):

P.E.M.D.A.S Order Of Operation

OpenStudy (mathstudent55):

You mean \(6 \times 3 \div 3 - 1 \) ? If so, this is how you do it step by step: \(6 \times 3 \div 3 - 1 \) \(= 18 \div 3 - 1\) \(= 6 - 1\) \( = 5\)